OpenStudy (anonymous):

please help Jill starts to save at age 25 for a vacation home that she wants to buy for her 50th birthday. She will contribute $2500 each year to an account, which earns 1.3% annual interest, compounded quarterly. What is the future value of this investment, rounded to the nearest dollar, when Jill is ready to purchase the vacation home? $73,712 $73,951 $294,847 $295,806

OpenStudy (anonymous):

@melissa_ @NaCl @mathmate @ivettef365

OpenStudy (anonymous):

A OR B let5 me check

OpenStudy (mathmate):

Use Accumulated amount = \( A\frac{R^n-1}{R-1} \) where A= amount contmibuted per period (of 3 months) = $2500/4 R=intereest per period of three months (0.013/4) n=number of periods (25 years * 4 quarters per year)